OpenStudy (anonymous):

Find x and RT if SU is a median of triangle RST if TU 2x+24 and UR 5x-30

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Since segments TU and UR are congruent, 2x + 24 = 5x - 30 Solve for x. Then RT = 2x + 24 + 5x - 30, or RT = 2(2x + 24) = 2(5x - 30)

OpenStudy (anonymous):

THANKK YOU :)

OpenStudy (mathstudent55):

You're welcome.